• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 12:04:16 on August 7,2005

    Entry number 151038

    keyword=Left Q1 reset.

    Mark Stevens and Gary Dezern have fixed the let arm Q1 problem.
    
    The reason that the interlocks would not reset is that the
    vacuum gauge measuring the magnet's insulating vacuum was reading back
    gibberish, and needed to be power cycled. After that it came
    back up reading nominal vacuum, and the interlocks could
    be reset. 
    
     Thanks Gary and Mark!